A force of 8.75N is applied to a mass of 3.5kg placed on a horizontal surface. If the coefficient of friction is placed between the mass and the surface is 0.2, what is the mass of acceleration?
1
Expert's answer
2011-10-14T12:05:35-0400
According to the Second Newton's Law:
ma = F - Friction.
Friction = 0.2 *
mg.
So,&
ma = F - 0.2mg. We get
a = F/m - 0.2g = 8.75/3.5 - 0.2*9.8 =
0.54 m/s^2